HALLAR EL CONJUNTO SOLUCION DE 2X^2-5X+2=X^2-2X+6
Answer by stanbon(75887)

2X^2-5X+2=X^2-2X+6
----------------------
x^2 - 3x -4 = 0
Factor:
(x-4)(x+1) = 0
x = 4 or x = -1
=======================
cheers,
Stan H.